OpenStudy (anonymous):

Solve the following for all real equations: 6x^2+6x=1

OpenStudy (campbell_st):

rewrite it as \[6x^2 + 6x + 1 = 0\] now just use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] and from your question a = 6, b = 5 and c = 1 hope it helps... the solutions are real but irrational

OpenStudy (anonymous):

im going to try it right now

OpenStudy (anonymous):

is the answer x= -1/2, -1/3?

OpenStudy (triciaal):

and from your question a = 6, b = 5 and c = 1 right formula b = 6

OpenStudy (triciaal):

a = 6, b = 6 c = -1

OpenStudy (anonymous):

aaaahhhhhhhhhhh how did you get that

OpenStudy (triciaal):

@campbell_st has the correct formula but the values for a, b and c he made an error. the equation is rewritten in the form ax^2 + bx + c =0 to pick up the values 6x^2+6x=1 6x^2 + 6x -1 = 0

OpenStudy (triciaal):

when I did it the values for x were -3 + rt 3; -3 -rt 3

OpenStudy (anonymous):

thank you

OpenStudy (triciaal):

you need to redo with c = -1

OpenStudy (triciaal):